# Token Is All You Need: Cognitive Planning through Sparse Intent Alignment

### v1

@@ Abstract
We challenge the long-standing assumption that exhaustive scene modeling is required for high-performance end-to-end autonomous driving (E2EAD). Unlike world-model approaches that rely on computationally intensive future scene generation or vision-language-action (VLA) systems constrained by Markov assumptions, we show that a minimal set of semantically rich tokens is sufficient for effective planning. Experiments on nuPlan (720 scenarios, 11k+ samples) with perception-informed BEV representations yield three key findings: (1) even without future prediction, our sparse representation achieves 0.548 m ADE—comparable to or surpassing prior methods reporting $\sim$0.75 m on nuScenes; (2) conditioning trajectory decoding on predicted future tokens reduces ADE to 0.479 m, a 12.6% improvement over current-state baselines; and (3) explicit reconstruction loss offers no benefit and may degrade performance under reliable perception inputs. Notably, we observe the emergence of temporal fuzziness: the model learns to attend adaptively to task-relevant semantics rather than aligning rigidly to fixed timestamps, providing a cognitive advantage for planning under uncertainty. Our “token is all you need” principle marks a paradigm shift—from reconstructing the world to understanding it—laying a foundation for cognitively inspired systems that plan through imagination rather than reaction.
@@ I Introduction
End-to-end autonomous driving (E2EAD) seeks to map perception directly to control, bypassing traditional pipelines that decompose perception, prediction, and planning [1, 2, 3]. Recent works such as SSR [4] achieve strong performance through sparse BEV representations and temporal self-supervision. Yet, most approaches still assume that effective planning requires dense scene representations and future self-supervision, leading to complex architectures that simulate full scene dynamics or rely on restrictive Markov assumptions.
World-model methods attempt to generate future scenes in latent space, but incur high computational cost and error accumulation in long-horizon reasoning. Likewise, vision-language-action (VLA) frameworks [5] introduce symbolic reasoning but typically assume temporal independence, limiting their ability to handle multi-modal or uncertain futures.
Humans, by contrast, neither reconstruct every visual detail nor act solely on the present. They maintain compact cognitive models, selectively attending to semantically relevant cues while mentally simulating possible futures. Inspired by this, we hypothesize that high-performance planning arises not from reconstructing the world but from understanding it—through a minimal set of semantically meaningful tokens encoding navigational intent.
We propose the Tokenized Intent World Model (TIWM), a cognitively inspired framework that redefines future reasoning in E2EAD:
@@
Experiments on nuPlan validate these principles. Even without future prediction, our sparse model achieves 0.548 m ADE, challenging the belief that dense modeling is essential. Conditioning trajectory decoding on predicted future tokens further improves ADE to 0.479 m—a 12.6% gain within the same framework—demonstrating that performance emerges from semantic imagination rather than pixel reconstruction. Adding explicit reconstruction loss provides no benefit, confirming that task-driven semantic alignment alone suffices under reliable perception.
Together, these results reveal a paradigm shift: effective planning stems not from exhaustive reconstruction but from selective understanding. By treating intent as the atomic unit of world representation, TIWM bridges two previously disjoint paradigms—world models that simulate and VLA systems that interpret—showing that belief–intent co-evolution, rather than dense reconstruction, underlies human-like planning.
@@ II Related Work
Autonomous driving has evolved from modular pipelines toward end-to-end (E2E) frameworks. Early systems decomposed perception, prediction, and planning into separate stages—interpretable but prone to error propagation and hand-crafted heuristics [6, 7]. Modern E2E approaches employ deep neural networks to map sensory inputs directly to control signals, enabling joint optimization across the full decision-making stack [1, 2, 3, 4].
@@
Vision-based methods such as UniAD [1] and VAD [2] achieve strong results via multi-camera fusion and multi-task learning. However, their reliance on dense feature reconstruction and short-horizon prediction limits long-term reasoning and scalability. In contrast, our work demonstrates that high-performance planning can emerge from a minimal set of semantically rich tokens—without dense reconstruction or auxiliary supervision.
@@
Latent world modeling frameworks such as World Models [8] and Dreamer [9] enable ”mental simulation” through compact latent dynamics. Yet, they face two major limitations in real-world driving: (1) the computational burden of generating full future scenes, and (2) the absence of theoretical guarantees for modeling multimodal behaviors, which leads to compounding errors in closed-loop control [10]. Our experiments show that this reconstruction is unnecessary—aligning sparse, task-relevant semantics with future intents suffices for high-quality planning. Unlike these scene-generative approaches, we focus solely on decision-critical information, showing that “token is all you need” for world modeling.
@@
Vision-language-action (VLA) frameworks unify perception, reasoning, and control [5], often conditioning on natural-language commands to enhance interpretability. However, most assume temporal independence between observations, producing reactive rather than anticipatory policies.
@@
Sparse scene modeling offers an alternative. SSR [4] extracts 16 navigation-guided tokens from BEV features and achieves 0.75 m ADE on nuScenes, using a Future Feature Predictor to align predicted and real BEV features via L2 loss. While conceptually aligned with our approach, SSR still relies on dense BEV feature reconstruction for self-supervision. As its authors note, “we supervise directly using an L2 loss with the real future BEV feature,” an assumption our results refute. We show that reconstruction loss not only fails to improve performance but can impede the emergence of high-level intent representations, whereas task-driven semantic alignment yields superior results.
@@
Most E2E planners—whether perception-to-trajectory regressors or VLM/VLA hybrids [11, 12]—remain constrained by Markovian dynamics, neglecting temporal dependencies and structural similarities in high-dimensional state spaces. This often leads to multi-objective conflicts and degraded safety generalization [10]. Consequently, such systems lack genuine “mental simulation,” i.e., the capacity to anticipate multiple futures before acting. Our method introduces autoregressive intent chains that enable future-oriented reasoning through task-driven semantic alignment rather than pixel-perfect reconstruction, capturing the essence of cognitive planning while avoiding over-constrained temporal alignment.
@@
Drawing from cognitive science, prior studies have explored attention, memory, and imagination in autonomous decision-making [13, 14, 15]. However, most remain conceptual or limited to architectural analogies. We instead operationalize two core cognitive principles—mental simulation [16, 17] and sparse coding [18]—within a unified generative–planning framework. This integration yields a computationally efficient yet cognitively grounded planner that achieves state-of-the-art performance.
@@ III Method
Figure 1: Tokenized Intent World Model: from perception to cognitive world. At each timestep, sparse tokens are extracted from BEV perception features, representing the distilled semantics of the scene. The model then autoregressively predicts future intent tokens—compact, task-relevant abstractions of the agent’s imagined goals. The planning decoder jointly reasons over current sparse tokens and predicted future intents to generate executable trajectories. Each intent token serves three roles: (1) Representation — compressing the core cognitive state of the environment; (2) Generation — unfolding plausible futures through autoregressive prediction; (3) Planning — directly guiding the trajectory decoder toward goal-directed action. In this framework, the world becomes an actionable cognitive entity—interpreted, imagined, and utilized by the decision system—rather than a dense intermediate reconstruction.
We propose the Tokenized Intent World Model (TIWM), a cognitively inspired planning framework that departs from conventional end-to-end autonomous driving (E2EAD) pipelines (Fig. 1). Unlike prior methods that rely on dense feature maps or pixel-level future reconstruction, TIWM achieves high-performance planning with a compact set of semantically rich tokens, requiring neither full-scene generation nor heavy self-supervision.
@@ III-A Sparse Intent Representation
TIWM operates on perception-informed Bird’s Eye View (BEV) representations, structured as a grid with two channels for dynamic objects and four for map layers (LANE, INTERSECTION, STOP_LINE, CROSSWALK). Instead of raw image inputs requiring depth estimation, TIWM leverages reliable perception modules that provide geometric and semantic stability suitable for real-world deployment. Since our goal is to study planning rather than perception, we assume access to reliable BEV inputs to isolate decision-level behavior.
@@
where attention weights are computed via a lightweight CNN:
@@
This process condenses a $256{\times}256$ BEV grid into 16 high-level semantic tokens, each representing a critical aspect of the driving scene. This follows the cognitive principle of sparse coding, where minimal active units capture the most relevant information for decision-making.
III-B Intent Chain Evolution and Task-Driven Alignment
Instead of explicitly predicting future states, TIWM models the evolution of intent chains through an autoregressive Transformer decoder:
$\mathbf{I}_{t+1}=\text{TransformerDecoder}(\mathbf{Q},\mathbf{T}_{t}),$ (3)
where $\mathbf{T}_{t}$ are current scene tokens, $\mathbf{Q}$ are learnable queries, and $\mathbf{I}_{t+1}$ is the predicted future intent. The decoder consists of two layers with eight attention heads and an embedding dimension of D=256D{=}256.
In contrast to SSR [4], which enforces dense alignment between predicted and real BEV features, TIWM learns task-driven semantic alignment. The training objective focuses purely on the planning outcome:
@@ III-B Intent Chain Evolution and Task-Driven Alignment
$\mathcal{L}_{\text{traj}}=\lambda_{\text{traj}}\|\text{TrajectoryPredictor}(\mathbf{T}_{t},\mathbf{I}_{t+1})-\mathbf{Y}_{t+1}\|_{2},$ (4)
and the overall objective is simply:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{traj}}.$ (5)
This minimalist formulation allows the model to adaptively focus on decision-critical semantics instead of rigid temporal matching. Empirically, it yields temporal fuzziness—an emergent behavior where the model flexibly attends to semantically meaningful elements across time, improving robustness under uncertainty.
For comparison, we also evaluate an auxiliary reconstruction loss:
@@
$\mathcal{L}_{\text{intent}}=\lambda_{\text{intent}}\|\mathbf{I}_{t+1}-\text{SparseExtractor}(\mathbf{Y}_{t+1})\|_{2},$ (6)
with a combined objective:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{traj}}+\mathcal{L}_{\text{intent}}.$ (7)
However, this explicit alignment consistently degraded performance by over-regularizing the token space, confirming that trajectory supervision alone is sufficient to learn meaningful intent representations. This supports our core principle: “token is all you need”—semantic alignment, not reconstruction, drives effective planning.
The resulting flexibility suggests that cognitive planning benefits from probabilistic, rather than deterministic, representations of future semantics. While this work focuses on a single intent chain, extending the model to multi-intent ensembles for representing alternative futures remains an exciting future direction.
III-C Cognitive Planning Mechanism
The trajectory decoder generates motion plans conditioned on both current scene tokens and predicted future intents:
$\mathbf{Y}_{t+1:t+H}=\text{TrajectoryPredictor}(\mathbf{T}_{t},\mathbf{I}_{t+1}),$ (8)
where $\mathbf{Y}_{t+1:t+H}$ denotes a 3-second trajectory over H=6H{=}6 future steps. The Multi-Modal Trajectory Predictor operates as follows:
  • Mode Selection: Predict mixture weights across $M$ driving modes based on the intent embedding:
    $\mathbf{w}=\text{softmax}(\mathbf{W}_{m}\mathbf{I}_{t+1}).$ (9)
  • Cross-Attention Fusion: Fuse current context and intent via multi-head attention:
    $\mathbf{Q}_{\text{fused}}=\text{LayerNorm}(\text{MultiHeadAttn}(\mathbf{Q},\mathbf{T}_{t},\mathbf{I}_{t+1})+\mathbf{Q}).$ (10)
  • Trajectory Generation: Decode fused queries into 2D trajectory points:
    $\mathbf{Y}=\text{MLP}(\mathbf{Q}_{\text{fused}}).$ (11)
This process embodies the cognitive notion of planning through imagination. Rather than reacting to the current state, the model anticipates possible futures through predicted intent tokens, evaluates alternatives, and plans actions accordingly. This yields a cognitively interpretable, foresight-driven decision mechanism. Empirically, removing explicit reconstruction loss ($\lambda_{\text{align}}=0$) improves both stability and accuracy, validating that task-driven semantic alignment alone enables robust planning—further reinforcing our guiding principle: “token is all you need.”
@@ IV-A Experimental Setup
Evaluation Metrics. Performance is measured by Average Displacement Error (ADE) and Final Displacement Error (FDE) over a 3 s horizon at 0.5 s intervals, following the standard AVG protocol used in VAD and SSR. ADE reflects mean trajectory deviation, while FDE captures the terminal-point error.
Implementation Details. TIWM adopts a ResNet-18 encoder (d=256d{=}256) and processes four historical frames (2 s) to predict six future points (3 s). Training uses AdamW ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$), batch size 32, learning rate $1{\times}10^{-4}$ with cosine decay, and weight decay $1{\times}10^{-4}$. All models are trained for 50 epochs on a single NVIDIA RTX 4090 (24 GB).

### v2

# Token Is All You Need: Cognitive Planning through Belief–Intent Co-Evolution

@@ Abstract
We challenge the long-standing assumption that exhaustive scene modeling is required for high-performance end-to-end autonomous driving (E2EAD). Inspired by cognitive science, we propose that effective planning arises not from reconstructing the world, but from the co-evolution of belief and intent within a minimal set of semantically rich tokens. Experiments on the nuPlan benchmark (720 scenarios, 11k+ samples) reveal three principles: (1) sparse intent tokens alone achieve 0.487 m ADE, demonstrating strong performance without future prediction; (2) conditioning trajectory decoding on predicted future tokens reduces ADE to 0.382 m, a 21.6% improvement, showing that performance emerges from cognitive planning; and (3) explicit reconstruction loss degrades performance, confirming that task-driven belief–intent co-evolution suffices under reliable perception inputs. Crucially, we observe the emergence of cognitive consistency: through prolonged training, the model spontaneously develops stable token dynamics that balance current perception (belief) and future goals (intent). This process, accompanied by ”temporal fuzziness,” enables robustness under uncertainty and continuous self-optimization. Our work establishes a new paradigm: intelligence lies not in pixel fidelity, but in the tokenized duality of belief and intent. By reframing planning as understanding rather than reaction, TIWM bridges the gap between world models and VLA systems, paving the way for foresightful agents that plan through imagination. Note: Numerical comparisons with methods reporting results on nuScenes are indicative only, as nuPlan presents a more challenging planning-focused evaluation.
@@ I Introduction
End-to-end autonomous driving (E2EAD) seeks to map perception directly to control, bypassing traditional pipelines that decompose perception, prediction, and planning [huPlanningOrientedAutonomousDriving2023a, jiangVADVectorizedScene2023a, zhengGenADGenerativeEndtoEnd2024]. Recent works such as SSR [liNavigationGuidedSparseScene2025] achieve strong performance through sparse BEV representations and temporal self-supervision. Yet, most approaches still assume that effective planning requires dense scene representations and future self-supervision, leading to complex architectures that simulate full scene dynamics or rely on restrictive Markov assumptions. World-model methods attempt to generate future scenes in latent space, but incur high computational cost and error accumulation in long-horizon reasoning. Likewise, vision-language-action (VLA) frameworks [zhouOpenDriveVLAEndtoendAutonomous2025] introduce symbolic reasoning but typically assume temporal independence, limiting their ability to handle multi-modal or uncertain futures. Humans, by contrast, neither reconstruct every visual detail nor act solely on the present. They maintain compact cognitive models, selectively attending to semantically relevant cues while mentally simulating possible futures. Inspired by this, we hypothesize that high-performance planning arises not from reconstructing the world but from understanding it—through a minimal set of semantically meaningful tokens encoding navigational intent. We propose the Tokenized Intent World Model (TIWM), a cognitively inspired framework that redefines future reasoning in E2EAD:
@@
  • Cognitive consistency learning: the system achieves persistent performance gains through prolonged training by dynamically balancing belief (current state understanding) and intent (future goal imagination), forming a self-organizing internal world model—a computational manifestation of belief–intent co-evolution.
Experiments on nuPlan validate these principles. Even without future prediction, our sparse model achieves 0.487 m ADE, challenging the belief that dense modeling is essential. Conditioning trajectory decoding on predicted future tokens further improves ADE to 0.382 m—a 21.6% gain within the same framework—demonstrating that performance emerges from semantic imagination rather than pixel reconstruction. Adding explicit reconstruction loss provides no benefit, confirming that task-driven semantic alignment alone suffices under reliable perception. Together, these results reveal a paradigm shift: effective planning stems not from exhaustive reconstruction but from selective understanding. By treating intent as the atomic unit of world representation, TIWM bridges two previously disjoint paradigms—world models that simulate and VLA systems that interpret—showing that belief–intent co-evolution, rather than dense reconstruction, underlies human-like planning.
@@ II Related Work
Autonomous driving has evolved from modular pipelines toward end-to-end (E2E) frameworks. Early systems decomposed perception, prediction, and planning into separate stages—interpretable but prone to error propagation and hand-crafted heuristics [liBEVFormerLearningBirdsEyeView2022, shiMTRMultiAgentMotion2024]. Modern E2E approaches employ deep neural networks to map sensory inputs directly to control signals, enabling joint optimization across the full decision-making stack [huPlanningOrientedAutonomousDriving2023a, jiangVADVectorizedScene2023a, zhengGenADGenerativeEndtoEnd2024, liNavigationGuidedSparseScene2025].
@@
Vision-based methods such as UniAD [huPlanningOrientedAutonomousDriving2023a] and VAD [jiangVADVectorizedScene2023a] achieve strong results via multi-camera fusion and multi-task learning. However, their reliance on dense feature reconstruction and short-horizon prediction limits long-term reasoning and scalability. In contrast, our work demonstrates that high-performance planning can emerge from a minimal set of semantically rich tokens—without dense reconstruction or auxiliary supervision.
@@
Latent world modeling frameworks such as World Models [haRecurrentWorldModels2018] and Dreamer [hafnerDreamControlLearning2020] enable ”mental simulation” through compact latent dynamics. Yet, they face two major limitations in real-world driving: (1) the computational burden of generating full future scenes, and (2) the absence of theoretical guarantees for modeling multimodal behaviors, which leads to compounding errors in closed-loop control [zhengSafeOfflineReinforcement2024]. Our experiments show that this reconstruction is unnecessary—aligning sparse, task-relevant semantics with future intents suffices for high-quality planning. Unlike these scene-generative approaches, we focus solely on decision-critical information, showing that “token is all you need” for world modeling.
@@
Vision-language-action (VLA) frameworks unify perception, reasoning, and control [zhouOpenDriveVLAEndtoendAutonomous2025], often conditioning on natural-language commands to enhance interpretability. However, most assume temporal independence between observations, producing reactive rather than anticipatory policies.
@@
Sparse scene modeling offers an alternative. SSR [liNavigationGuidedSparseScene2025] extracts 16 navigation-guided tokens from BEV features and achieves 0.75 m ADE on nuScenes, using a Future Feature Predictor to align predicted and real BEV features via L2 loss. While conceptually aligned with our approach, SSR still relies on dense BEV feature reconstruction for self-supervision. As its authors note, “we supervise directly using an L2 loss with the real future BEV feature,” an assumption our results refute. We show that reconstruction loss not only fails to improve performance but can impede the emergence of high-level intent representations, whereas task-driven semantic alignment yields superior results.
@@
Most E2E planners—whether perception-to-trajectory regressors or VLM/VLA hybrids [tianDriveVLMConvergenceAutonomous2024, bansalChauffeurNetLearningDrive2018]—remain constrained by Markovian dynamics, neglecting temporal dependencies and structural similarities in high-dimensional state spaces. This often leads to multi-objective conflicts and degraded safety generalization [zhengSafeOfflineReinforcement2024]. Consequently, such systems lack genuine “mental simulation,” i.e., the capacity to anticipate multiple futures before acting. Our method introduces autoregressive intent chains that enable future-oriented reasoning through task-driven semantic alignment rather than pixel-perfect reconstruction, capturing the essence of cognitive planning while avoiding over-constrained temporal alignment.
@@
Drawing from cognitive science, prior studies have explored attention, memory, and imagination in autonomous decision-making [dalioExploitingDreamlikeSimulation2017, mahmoudSelfdrivingCarsLearn2018, mazzalaiCognitivelyInspiredFramework2018]. However, most remain conceptual or limited to architectural analogies. We instead operationalize two core cognitive principles—mental simulation [decetyPowerSimulationImagining2006, mattarPlanningBrain2022] and sparse coding [olshausenEmergenceSimplecellReceptive1996]—within a unified generative–planning framework. This integration yields a computationally efficient yet cognitively grounded planner that achieves state-of-the-art performance.
@@ III Method
Figure 1: Tokenized Intent World Model: from perception to cognitive world. At each timestep, sparse tokens are extracted from BEV perception features, representing the distilled semantics of the scene. The model then autoregressively predicts future intent tokens—compact, task-relevant abstractions of the agent’s imagined goals. The planning decoder jointly reasons over current sparse tokens and predicted future intents to generate executable trajectories. Each intent token serves three roles: (1) Representation — compressing the core cognitive state of the environment; (2) Intent — unfolding plausible futures through autoregressive prediction; (3) Decision — directly guiding the trajectory decoder toward goal-directed action. In this framework, the world becomes an actionable cognitive entity—interpreted, imagined, and utilized by the decision system—rather than a dense intermediate reconstruction.
We propose the Tokenized Intent World Model (TIWM), a cognitively grounded planning framework that redefines end-to-end autonomous driving (E2EAD) as a process of belief–intent co-evolution (Fig. 1). Unlike conventional approaches that rely on dense feature maps or pixel-level future reconstruction, TIWM achieves high-performance planning through a compact set of semantically rich tokens, requiring neither full-scene generation nor auxiliary self-supervision.
@@ III-A Sparse Intent Representation
TIWM operates on perception-informed Bird’s Eye View (BEV) representations, structured as a grid with two channels for dynamic objects and four for semantic map layers (LANE, INTERSECTION, STOP_LINE, CROSSWALK). Following standard practice, we assume access to reliable BEV inputs from a perception stack, thereby isolating the study of decision-level behavior.
@@
where the attention weights are computed via a lightweight CNN:
@@
This process condenses a $256{\times}256$ BEV grid into 16 high-level semantic tokens, each representing a critical aspect of the driving scene. This design operationalizes the cognitive principle of sparse coding [olshausenEmergenceSimplecellReceptive1996], where minimal active units capture the most decision-relevant information.
III-B Intent Chain Evolution via Belief–Intent Co-Evolution
TIWM models planning as the dynamic interplay between belief (understanding of the present) and intent (imagination of the future). Instead of predicting full future states, the model autoregressively evolves an intent chain from a sequence of historical tokens.
Given four historical token sets $\mathbf{T}_{t-3:t}$, a learnable future query $\mathbf{Q}\in\mathbb{R}^{N\times D}$, and a causal attention mask $\mathcal{M}$, we compute the future intent $\mathbf{I}_{t+1}$ as:
$\begin{split}[\mathbf{T}_{t-3:t},\mathbf{I}_{t+1}]=&\text{TransformerEncoder}\\ &(\text{Concat}(\mathbf{T}_{t-3:t},\mathbf{Q});\mathcal{M}),\end{split}$ (3)
The TransformerEncoder consists of two layers with eight attention heads and an embedding dimension of $D=256$. The causal mask $\mathcal{M}$ ensures that future intent tokens can only attend to past and present information, preventing information leakage.
This architecture enables belief–intent co-evolution: the current tokens $\mathbf{T}_{t}$ form the belief state, while the predicted $\mathbf{I}_{t+1}$ constitutes the intent state. They are not independent; rather, they are jointly shaped by a single task-driven objective, leading to a coherent internal world model.
III-C Task-Driven Semantic Alignment and Cognitive Consistency
TIWM’s training objective is intentionally minimalist, relying solely on the final planning outcome:
@@ III-C Task-Driven Semantic Alignment and Cognitive Consistency
$\begin{split}\mathcal{L}_{\text{traj}}=\lambda_{\text{traj}}\cdot\text{SmoothL1}&(\text{TrajectoryPredictor}(\mathbf{T}_{t},\mathbf{I}_{t+1}),\\ &\mathbf{Y}),\end{split}$ (4)
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{traj}},$ (5)
where $\mathbf{Y}_{t+1:t+H}$ is the ground-truth 3-second trajectory ($H=6$).
For ablation, we also evaluate an auxiliary reconstruction loss that aligns the predicted intent with a ground-truth token:
@@
$\begin{split}\mathcal{L}_{\text{intent}}=\lambda_{\text{intent}}&\|\text{F.normalize}(\mathbf{I}_{t+1})\\ &-\text{F.normalize}(\mathbf{T}_{t+1}^{\text{gt}})\|_{2},\end{split}$ (6)
with $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{traj}}+\mathcal{L}_{\text{intent}}$. However, this explicit supervision consistently degrades performance, confirming that trajectory-driven learning alone is sufficient—and superior—for inducing meaningful intent representations.
Through prolonged training, this task-driven process leads to the emergence of cognitive consistency: the belief–intent pair self-organizes into a stable, robust equilibrium. This is not a static state but a dynamic balance, akin to a ”quantum state transition” between representing ”what is” and ”what could be,” enabling persistent self-optimization over thousands of epochs.
III-D Cognitive Planning Mechanism
The final trajectory is generated by a Multi-Modal Trajectory Predictor:
$\mathbf{Y}_{t+1:t+H}=\text{TrajectoryPredictor}(\mathbf{T}_{t},\mathbf{I}_{t+1}).$ (7)
The predictor first computes mode weights $\mathbf{w}=\text{softmax}(\mathbf{W}_{m}\mathbf{I}_{t+1})$, then fuses current context and intent via cross-attention, and finally decodes the fused representation into 2D trajectory points. This process embodies planning through imagination, where the agent reasons over possible futures before committing to an action.
@@ IV-A Experimental Setup
Evaluation Metrics. Performance is measured by Average Displacement Error (ADE) and Final Displacement Error (FDE) over a 3 s horizon at 0.5 s intervals, following the standard MAX protocol used in UniAD and SSR. ADE reflects mean trajectory deviation, while FDE captures the terminal-point error.
Implementation Details. TIWM adopts a ResNet-18 encoder (d=256d{=}256) and processes four historical frames (2 s) to predict six future points (3 s). Training uses AdamW ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$), batch size 32, learning rate $1{\times}10^{-4}$ with cosine decay, and weight decay $1{\times}10^{-4}$. All models are trained on a single NVIDIA RTX 4090 (24 GB).